# leftmargin=3.2em, rightmargin=3.2em

### v2

@@ Conclusion
Reliance drills are novel exercises that can help organisations to mitigate human over-reliance on AI systems in real-world settings. They can be implemented through a five-step pipeline that guides an organisation when designing the drill, evaluating its risks, and deciding on an appropriate response to over-reliance. Ultimately, these drills could become a valuable tool that helps safety-critical industries to harness the benefits of AI while guarding against the risk posed by over-reliance.
@@
Reliance drills are designed to mitigate human over-reliance on AI systems, and we hope that they will be adopted as a standard risk management practice. However, we also appreciate that these drills could introduce novel safety risks, negatively impact users’ mental health, or inadvertently induce under-reliance. Therefore, the paper directly addressed these concerns in Section 3.
@@ Acknowledgments
Appendix A. Incentive structures for reliance drills
This appendix discusses potential incentive structures for reliance drills. To frame this discussion, we explore the incentives for a widely-adopted predecessor of reliance drills: phishing simulations. During one of these simulations, a company’s employees are deliberately sent (fake) scam emails to determine whether they are able to recognise a phishing attack [24, 18]. In many ways, reliance drills generalise this approach, as they test whether a company’s employees can recognise a broad range of misleading (AI-generated) content, not just scam emails. Given the close similarity between these techniques, we explore whether the business incentives that facilitate phishing simulations could serve as a blueprint for reliance drills.
Large organisations often have the resources needed to conduct in-house phishing simulations, where an internal team creates and sends (fake) scam emails. This in-house approach can be particularly effective as an internal team can easily tailor their (fake) phishing campaign to address company-specific vulnerabilities. Alternatively, some companies might hire an external specialist to run phishing simulations for them. These external providers attack their client’s employees with a simulated phishing campaign and then write a safety report to summarise the results. This outsourced approach may be appropriate for organisations that lack the expertise to run their own simulations.
Notably, it may be possible to adapt the business model used by phishing simulations for reliance drills. Namely, large businesses may conduct their own in-house drills, while smaller organisations might hire third-party specialists. Crucially, however, for this business model to materialise, managers will need an incentive to pay for these drills. For instance, there is a clear incentive for phishing simulations as they can reduce the probability of a cyber-security breach, thereby lowering the cost of insurance premiums that cover this risk. In the remainder of this appendix, we identify and examine two possible incentives for reliance drills.
Reputational. Many organisations hesitate to adopt AI systems due to the reputational damage of an AI-induced accident [9]. If these organisations are to eventually adopt AI systems, they might require risk management strategies (such as reliance drills) that can help to allay their concerns.
Financial. When organisations deploy AI systems, they may purchase liability insurance that hedges against the risk of a costly accident [13]. This creates a financial incentive for conducting reliance drills, as insurers typically offer lower premiums to organisations with risk management practices. Additionally, if the providers of an AI system are held partially liable for an accident, they would also have a financial incentive to build reliance drills into their products. Ultimately, the extent and distribution of these financial incentives will depend on decisions by regulators and liability lawyers.
In summary, while the analogy between phishing simulations and reliance drills is imperfect, it provides a starting point to analyse potential incentive structures for reliance drills. For example, we identified that reliance drills might be run within an organisation or conducted by a third-party specialist. Building on this, we outlined potential incentives that might compel an organisation to pay for reliance drills, or something like them. We hope that future work will further explore the regulations, incentives, and market structures that might facilitate reliance drills.
@@ Appendix B. Empirical evaluation of reliance drill efficacy
This paper proposes the reliance drill as a tool for identifying human over-reliance on AI systems. In Section 3, we described various methods that an organisation might use to correct over-reliance, once it has been identified [16, 12]. However, suppose that an organisation wanted to gather their own evidence to assess whether reliance drills, coupled with a simple warning system, can measurably reduce users’ over-reliance on AI. This appendix outlines a simple experiment that they could conduct.
@@
    Group 3 - Reliance drill: Participants receive the same AI assistance as Group 2, but are subjected to a reliance drill early in the test. Specifically, for 1-2 non-rated questions, the AI’s advice would be modified so that it becomes incorrect. After answering one of these questions, participants would receive feedback on their performance in the reliance drill before continuing with the test. [^2] Additionally, researchers could also decide to add a fourth group that still receives the same 1-2 (deliberately mistaken) non-rated questions but does not receive feedback on their performance.
Interpreting the results. There are at least two key metrics in this experiment. First, a researcher could measure the level of over-reliance. To quantify this, they might count the number of questions that more participants answer correctly in Group 1 when compared with Group 2 and Group 3, respectively. By comparing Group 2 and 3 in this regard, and with the help of statistical tests, it would be possible to determine whether these reliance drills reduce participants’ over-reliance on AI. [^3] This analysis of the IIT (intent-to-treat-effects) could be complemented with an estimate of non-compliance in the experimental conditions, such as a self-report assessment of how much each user relied on the AI’s advice.
@@
Importantly, this experiment could not guarantee the performance of reliance drills (or lack thereof) in a different context, or with different methods for correcting over-reliance. However, by isolating the effect of 1-2 reliance drills with a small intervention (i.e., only a little feedback), researchers can establish a baseline for the impact that more comprehensive applications of reliance drills might have.
Appendix C. A broader taxonomy of human reliance on AI
While this paper focuses on over-reliance, we also recognise that there is a broader taxonomy of human reliance on AI systems, shown in Figure 3. Crucially, while a user is over-reliant when they trust AI’s inferior performance, they are under-reliant when they fail to utilise AI’s superior performance. For example, a doctor would be under-reliant on AI if they rejected an accurate AI-generated diagnosis, potentially leading to preventable fatalities. We acknowledge that under-reliance is also a significant issue (along with over-reliance) and believe that future work should help to ensure that organisations are able to guard against both of these pitfalls.
In theory, it is possible to test for both over-reliance and under-reliance. Figure 3 provides a basis for potential future experiments: One could either randomise column-wise (e.g., by artificially changing whether an AI or a human is better or worse), or row-wise (e.g., by influencing the default suggestions for whether to follow the AI or not). Ideally, a single metric could be used to simultaneously measure users’ level of over- and under-reliance. While some researchers have already started to develop such a metric [7], more research is needed to determine how this could be applied in real-world settings.
@@ Appendix D. Applying reliance drills in a military setting
This appendix expands on the approach taken in Section 4. In it, we outline a hypothetical military scenario where reliance drills could be used to reduce over-reliance on AI systems. Specifically, we examine how AI might shape the job of command and control operators that are responsible for detecting and responding to hostile military activity.
@@
Set criteria for impairing AI In a military operation, time is extremely valuable, but hasty decisions based on incorrect information can be fatal. Therefore, the military’s leadership might be interested in observing how operators react to a broad range of mistakes during a reliance drill. Some of the AI’s mistakes could be severe—with potentially lethal outcomes—while others might be relatively mild (such as slightly overemphasising a benign signal). Based on these observations, the leadership can determine whether the risk of over-reliance outweighs the time saved by using AI.
Perform risk assessment Given the level of risk, these drills must not be conducted in a real-world setting. Instead, they should be run in a dedicated training environment. Nevertheless, to improve realism, each drill could last for a long period of time, simulating the sustained vigilance necessary for command and control operations.
@@
During a reliance drill, an AI system would make a variety of errors. Operators should identify these mistakes by cross-referencing the AI’s conclusions with other available data. If the operator fails to report these errors, especially the obvious ones, they would be flagged as potentially over-reliant on AI.
@@
In a debrief, an investigator must check that operators have not become excessively sceptical of AI systems. Additionally, given the intense nature of these drills, the debrief should highlight mental health services that operators can contact for support.

### v3

@@ Conclusion
Reliance drills are a novel safety practice that organisations can use to mitigate human over-reliance on AI assistance. By following a five-step pipeline, organisations can design these drills, evaluate their risks, and decide on appropriate responses to instances of over-reliance. Ultimately, these drills could become a valuable tool for safety-critical industries, allowing them to harness the benefits of AI while guarding against the risks associated with over-reliance.
@@
We hope that reliance drills will be adopted as a standard risk management procedure. However, we also appreciate that they could introduce novel safety risks, negatively impact users’ mental health, or inadvertently induce under-reliance. This paper addresses each of these concerns in Section 4.
@@ Acknowledgments
Appendix A. Background literature on technological over-reliance
Technological over-reliance, where humans depend excessively on external tools and systems, has been extensively studied across various domains. [^3] Some researchers distinguish between individual over-reliance and systemic over-reliance on AI, where the latter considers how societally vital functions dependent on AI. Our paper is mostly focused on the former. This appendix provides a brief overview of the relevant terminology and situates our work within this broader context. Researchers have developed several overlapping concepts while studying over-reliance, some of which are outlined below:
  1. 1.
    Automation bias. Empirical studies have consistently found a psychological vulnerability, where humans place excessive trust in the decisions made by automated systems, particularly in situations that are complex or subject to time constraints [12, 37, 28].
  2. 2.
    Human-centred design. To mitigate over-reliance, some researchers advocate for computer systems that are designed to actively support user autonomy. For instance, computers could encourage their users to pause and reflect between human-computer interactions [4].
  3. 3.
    Explainable AI. One branch of over-reliance research is focused on ensuring that computers provide faithful and interpretable explanations for their decisions. In theory, if computers act transparently, then humans can understand and thereby scrutinise their decisions [41].
  4. 4.
    Human-in-the-loop. For critical decisions, many researchers argue that computers should not be allowed to act autonomously. Instead, they recommend mandatory human oversight, where humans authorise important decisions before they are taken [8, 14].
Each of these concepts has helped researchers better understand technological over-reliance. In medicine, for example, numerous studies have used these concepts to determine whether doctors are over-reliant on AI tools when diagnosing patients [1, 11, 20]. While these concepts have provided important insights about over-reliance in specific contexts, they have not been translated into a standardised test that organisations can use to monitor over-reliance. Our paper has argued for just that, offering reliance drills as a tractable risk management strategy that can be used to monitor and mitigate humans’ over-reliance on AI systems. [^4] Reliance drills do not measure over-reliance as defined in the broadest sense, which covers every way that a technology could negatively affect humans’ performance and autonomy. Instead, reliance drills measure a small, tractable component of this phenomenon: whether users can identify mistakes in supposedly AI-generated text.
@@ Appendix B. Empirical evaluation of reliance drill efficacy
This paper proposes reliance drills as an approach to identify and mitigate human over-reliance on AI systems. However, before spending time and resources to implement these drills, organisations may want evidence that they are worth the investment. Admittedly, in certain situations, where over-reliance is not of particular concern, reliance drills may not be worth the cost. However, in other scenarios, especially where instances of over-reliance could result in large financial liabilities, reliance drills could be an incredibly valuable safety practice.
How might an organisation determine whether reliance drills are useful for their specific application? To perform an informed cost-benefits analysis, organisations may choose to gather their own evidence about these drills. For example, they could run a simple experiment that assesses whether reliance drills, coupled with a simple warning system, can measurably reduce users’ over-reliance on AI. This appendix outlines how such an experiment could be implemented.
@@
    Group 3 - Reliance drill: Participants receive the same AI assistance as Group 2, but are subjected to a reliance drill. Specifically, for 1-2 non-rated questions, the AI’s advice would be modified to be incorrect. Upon answering these questions, participants would receive feedback on their performance in the reliance drill before continuing with the test. [^5] Researchers could also add a fourth group that receives the same 1-2 (deliberately mistaken) non-rated questions but does not receive feedback on their performance.
Interpreting the results. There are at least two key metrics in this experiment. First, a researcher could measure the level of over-reliance. To quantify this, they might count the number of questions that more participants answered correctly in Group 1 when compared with Group 2 and Group 3, respectively. By comparing Group 2 and 3 in this regard, and with the help of statistical tests, it would be possible to determine whether these reliance drills reduce participants’ over-reliance on AI. [^6] This analysis of the IIT (intent-to-treat-effects) could be complemented with an estimate of non-compliance in the experimental conditions, such as a self-report assessment of how much each user relied on the AI’s advice.
@@
Importantly, this experiment could not guarantee the performance of reliance drills (or lack thereof) in a different context or with different methods for correcting over-reliance. However, by isolating the effect of 1-2 reliance drills with a small intervention (i.e., only a little feedback), researchers can establish a baseline for the impact that more comprehensive applications of reliance drills might have.
Appendix C. A broader taxonomy for human reliance on AI
While this paper focuses on over-reliance, we also recognise that there is a broader taxonomy for human reliance on AI systems, shown in Figure 3. Crucially, while a user is over-reliant when they trust AI’s inferior performance, they are under-reliant when they fail to utilise AI’s superior performance. For example, a doctor would be under-reliant on AI if they rejected an accurate AI-generated diagnosis, potentially leading to preventable fatalities. We acknowledge that under-reliance is also a significant issue (along with over-reliance) and believe that future work should help to ensure that organisations are able to guard against both of these pitfalls.
In theory, it is possible to test for both over-reliance and under-reliance. Figure 3 provides a basis for potential future experiments: One could either randomise column-wise (e.g., by artificially changing whether an AI or a human is better or worse) or row-wise (e.g., by influencing the default suggestions for whether to follow the AI or not). Ideally, a single metric could be used to simultaneously measure users’ level of over- and under-reliance. While some researchers have already started to develop such a metric [15], more research is needed to determine how this could be applied in real-world settings.
@@ Appendix D. Applying reliance drills in a military setting
This appendix expands on the approach taken in Section 5. In it, we outline a hypothetical military scenario where reliance drills could be used to reduce over-reliance on AI. We examine how AI might impact command and control operators that detect and respond to hostile military activity.
@@
Decide how to
impair the AI
In a military operation, time is extremely valuable, but hasty decisions based on incorrect information can be fatal. Therefore, the military’s leadership might be interested in observing how operators react to a broad range of supposedly AI-generated mistakes during a reliance drill. Some of the AI system’s mistakes could be severe—with potentially lethal outcomes—while others might be relatively mild (such as slightly overemphasising a benign signal). Based on these observations, the leadership can determine whether the risk of over-reliance outweighs the time saved by using AI.
Perform risk
assessment
Given the level of risk, these drills must not be conducted in a real-world setting. Instead, they should be run in a dedicated training environment. Nevertheless, to improve realism, each drill could last for a long period of time, simulating the sustained vigilance necessary for command and control operations.
@@
During a reliance drill, an AI system would make a variety of errors. Operators should identify these mistakes by cross-referencing the AI system’s conclusions with other available data. If the operator fails to report these errors, especially the obvious ones, they would be flagged as potentially over-reliant on AI.
@@
Given the intense nature of these drills, during a debrief, investigators should highlight mental health services that operators can contact for support.